Identify the appropriate distribution function if we want to select 20 random numbers between 0 and 9.




1
Expert's answer
2022-06-09T12:07:08-0400

Uniform Distribution is appropriate distribution.


The random-number generator on calculators randomly generates a number between 0 and 9. The random variable "X," the number generated, follows a uniform probability distribution.

The notation for the uniform distribution is "X{\sim} U(a, b)" where "a=" the lowest value of "x" and "b=" the highest value of "x."

The probability density function is


"f(x)=\frac{1}{b-a}"


for "a \leq x \leq b."

Formulas for the theoretical mean and standard deviation are


"\mu=\dfrac{a+b}2,\sigma=\sqrt{\dfrac{(b-a)^2}{12}}"



"a=0, b=9"

The probability density function is


"f(x)=\frac{1}{9-0}"


for "0 \leq x \leq 9."

Formulas for the theoretical mean and standard deviation are


"\mu=\dfrac{0+9}2=4.5,\sigma=\sqrt{\dfrac{(9-0)^2}{12}}=0.25"